Determine the measure of each angle of a regular pentagon, regular nonagon.


1
Expert's answer
2021-01-07T18:03:17-0500

Regular pentagon





"\alpha=\dfrac{360\degree}{5}=72\degree"




"\beta=\dfrac{180\degree(5-2)}{5}=108\degree"




"interior\ angle=108\degree"

The interior and exterior angle add up to 180°


"exterior\ angle=180\degree-\beta=180\degree-108\degree"

"exterior\ angle=72\degree"


Regular nonagon





"\alpha=\dfrac{360\degree}{9}=40\degree"




"\beta=\dfrac{180\degree(9-2)}{9}=140\degree"




"interior\ angle=140\degree"

The interior and exterior angle add up to 180°


"exterior\ angle=180\degree-\beta=180\degree-140\degree"

"exterior\ angle=40\degree"